OpenStudy (anonymous):

Given RS = 6, ST = 3, and SL = 4, find the required measure.

OpenStudy (anonymous):

Using Pythagoras theorem, SM =Sqrt(3*3+4*4) = Sqrt(9+16) = Sqrt(25) = 5.

OpenStudy (anonymous):

im on odyssey ware and its saying something about a six?

OpenStudy (anonymous):

im not the best at math so dont trust my answer 100%♫♪

OpenStudy (anonymous):

lol me either thats why i put it on here

OpenStudy (anonymous):

im sorry all i know to do is keep trying ♥